# Multi-modal Stance Detection: New Datasets and Model

### v2

@@ 6.2 Zero-shot Multi-modal Stance Detection
The results of zero-shot multi-modal stance detection are reported in Table 5. It can be seen that the large language models achieve superior performance due to the need for detecting stances on unseen targets. This may be attributed to the powerful zero-shot learning ability of large language models For our TMPT, which does not use chain-of-thoughts from large language models, still achieves better performance than large language models in some targets, while also outperforming all non-large language model baselines. This shows the promise of our TMPT in multi-modal stance detection. Further, TMPT+CoT performs overall better than TMPT, which indicates that obtaining powerful text and visual comprehension abilities from large models may be key to improving the performance of detecting stance for unseen targets.
@@
w/o $\mathcal{P}^{V}$ 64.86 69.66 78.06 59.15 71.73
Table 6: Experimental results of ablation study. The reported results are the average results of all targets in a dataset on in-target multi-modal stance detection.
@@ 6.3 Ablation Study
To analyze the impact of the targeted prompt tuning in our proposed TMPT, we conduct an ablation study and report the results in Table 6. Note that the removal of textual prompting (w/o $\mathcal{P}^{T}$) sharply degrades the performance, which verifies the significance of textual prompting in learning targeted stance features for multi-modal stance detection. In addition, the removal of visual prompt tuning (w/o $\mathcal{P}^{V}$) leads to considerable performance degradation, which indicates that utilizing visual prompt tuning can make better learning of multi-modal targeted stance information and thus improves the performance of multi-modal stance detection.
Figure 3: Performance of using different pre-trained language models: Roberta Nguyen et al. (2020) (left) and KEBERT Kawintiranon and Singh (2021) (right). The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Figure 4: Performance of using different multi-modal fusion methods. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
@@ 6.4 Generalization of Targeted Multi-modal Prompt Tuning
Previous experiments have demonstrated that the stance expression primarily resides in the textual modality. Therefore, to investigate the generalization of our TMPT when used with different pre-trained language models, we conduct experiments with two variants of our TMPT by using two other promising PLMs: RoBERTa Liu et al. (2019b) and KEBERT Kawintiranon and Singh (2021). The results are shown in Figure 3. Given the page constraints, we conduct experiments on Mtse, Mccq, and Mwtwt datasets. Note that our Targeted Multi-modal Prompting can directly work with the two PLMs and achieve better performance, showing the compatibility of our method with various pre-trained models.
@@
We choose cross-modal attention Wei et al. (2020) and GAT Velickovic et al. (2018) to analyze the performance of our TMPT with different multi-modal fusion methods. Given the page constraints, we conducted our experiments on Mtse, Mccq, and Mwtwt datasets. The results are shown in Figure 4. It can be seen that no matter which fusion method we use, the experimental performance is better than BERT+ViT. This indicates that our TMPT can directly work with various multi-modal fusion methods and lead to improved multi-modal stance detection performance.
@@ 7 Conclusion
This paper presents Mtse, Mccq, Mwtwt, Mruc and Mtwq, five new datasets for multi-modal stance detection. Based on the created datasets, we present in-target multi-modal stance detection and zero-shot multi-modal stance detection, aiming to advance and facilitate research in the field of multi-modal stance detection. In addition, we propose a simple yet effective targeted multi-modal prompting framework (TMPT) to deal with multi-modal stance detection, where the target information is explored to prompt the pre-trained models in learning multi-modal stance features. Extensive experiments on the new datasets show that our proposed TMPT achieves overall better performance than state-of-the-art baseline methods.
@@ A.1 Multi-modal Twitter Stance Election 2020
    Trump 01/01/2020 $\rightarrow$ 09/30/2020 Biden 01/01/2020 $\rightarrow$ 09/30/2020
@@ A.4 Multi-modal Russo-Ukrainian Conflict
    Russia 01/01/2022 $\rightarrow$ 06/30/2023 Ukraine 01/01/2022 $\rightarrow$ 06/30/2023
@@ A.5 Multi-modal Taiwan Question
    Mainland of China 01/01/2022 $\rightarrow$ 06/30/2023 Taiwan of China 01/01/2022 $\rightarrow$ 06/30/2023
@@ C.1 Analysis of Textual Prompt Tuning
For textual prompt tuning, we utilize a frozen paradigm to make full use of the semantic information learned by the pre-trained language model. To analyze the effectiveness of the frozen paradigm, we also tried a tuned prompt, the results are shown in Table 7. The results of the tuned paradigm are extremely poorer than the frozen one and fluctuate greatly. One possible reason is that, unlike the continuous picture patch in the visual transformer, the token in the textual modality is discrete single words, so it is hard to use a gradient-based method to fine-tune the prompts like a visual modality. Thus, in this paper, we choose a manually designed fixed prompt to obtain better performance.
@@
Table 7: Experimental results of fixed prompt and tuned prompt of textual modality. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
@@
To analyze the impact of the type of the hand-design textual prompts, we design several types of textual prompts and report the experiments in Table 8. Take the MTSE2020 dataset as an example, it can be seen that the experimental results of all types of prompts are superior to those without prompts (BERT+ViT). This demonstrates the significance of the targeted prompts in this task. Further, we can also see that there are considerable differences in performance between different types of prompts. Specifically, when using both target and stance to design prompts (③, ④, and ⑤), the model performs significantly better. Therefore, in our method, we choose the type of ⑤ to devise the textual prompts for multi-modal stance detection.
@@
BERT+ViT - 63.28
Trump 63.32
Donald Trump 63.57
stance on Donald Trump 64.67
What is the stance on Donald Trump? 64.59
@@ C.2 Analysis of Visual Prompt Tuning
Following Jia et al. (2022), we conduct experiments with different depths of visual prompt tuning to analyze the impact of the depth of prompt tuning in our model. Given the page constraints, we conducted our experiments on MTSE, MCCQ, and MWTWT. The experimental results are shown in Table 9. We can see that both shallow and deep prompt tuning are good in our model, while the results of shallow prompt tuning are slightly better. Therefore, we adopt shallow prompt tuning in our method.
Method Mtse Mccq Mwtwt
Shallow (ours) 66.12 70.03 79.03
@@
Table 9: Comparison results of using the shallow prompt tuning and deep prompt tuning. Shallow prompt tuning refers to only fine-tuning the prompt tokens in the Embedding layer of ViT, while deep prompt tuning refers to fine-tuning the prompt tokens in the Embedding layer of ViT and each layer in the Transformer. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
@@
To analyze the impact of the number of visual prompt tokens, we set the value range of tokens as {3, 5, 7, 9} for comparative experiments. Given the page constraints, we conducted our experiments on MTSE, MCCQ, and MWTWT. The results are shown in Figure 6. Note that different values of tokens can have a certain impact on performance. When the value is 7, the overall performance of the model on all datasets is the best. Therefore, we set the number of tokens to 7 in our method.
Figure 6: Experimental results of using different numbers of visual prompting tuning tokens. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Initialization of prompting tuning tokens
To analyze the impact of the initialization of prompting tuning tokens, similar to textual modality, we use target-related images to initialize prompt tokens. For example, on the Mtse dataset, we use the official portraits of two targets to initialize the corresponding visual prompting tuning tokens. The results are shown in Table 10. We can see that initializing prompting tuning tokens with images related to the target can achieve better results than random initialization.
Method Mtse
Initializing with related pictures (ours) 66.66
[…]
Table 10: Experimental results of using target-related images to initialize prompt tokens (ours) and random initialization. The reported results are the average results of all targets in one dataset on in-target multi-modal stance detection.
Appendix D Analysis of Our Datasets and Task
By comparing the results of BERT and ViT in Table 4 and 5, we ascertain that in the five multi-modal stance detection datasets, text modality is the dominant modality, comprising more stance information. Further, as described in Section 3.4 the annotators determine that nearly half of the samples contain stance information in their images, which maintains that the visual modality is important to facilitate textual modality in detecting stance, which can also be verified by the experimental results in Table 4 and 5 that BERT+ViT outperforms BERT and ViT and the visualization in Figure 5. In addition, from the results in Table 4 and 5, we can also see that even though our model performs unsatisfactorily on most datasets in all stance detection scenarios, which validates the challenge of our datasets and this task from another perspective.

### v3

@@ 6.2 Zero-shot Multi-modal Stance Detection
The results of zero-shot multi-modal stance detection are reported in Table 6. It can be seen that the large language models achieve superior performance due to the need for detecting stances on unseen targets. This may be attributed to the powerful zero-shot learning ability of large language models. For our TMPT, which does not use chain-of-thoughts from large language models, still achieves better performance than large language models in some targets, while also outperforming most of non-large language model baselines. This shows the promise of our TMPT in zero-shot stance detection. Further, TMPT+CoT performs overall better than TMPT, which indicates that obtaining powerful text and visual comprehension abilities from large models may be key to improving the performance of detecting stance for unseen targets.
@@
w/o $\mathcal{P}^{V}$ 58.14 65.71 73.93 70.85 63.69
Table 7: Experimental results of ablation study. The reported results are the Macro F1-score across all targets in a dataset on in-target multi-modal stance detection.
@@ 6.3 Ablation Study
To analyze the impact of the targeted prompt tuning in our proposed TMPT, we conduct an ablation study and report the results in Table 7. Note that the removal of textual prompting (w/o $\mathcal{P}^{T}$) sharply degrades the performance, which verifies the significance of textual prompting in learning textual targeted stance features for multi-modal stance detection. In addition, the removal of visual prompt tuning (w/o $\mathcal{P}^{V}$) leads to considerable performance degradation, which indicates that utilizing visual prompt tuning can make better learning of visual targeted stance information and thus improves the performance of multi-modal stance detection.
Figure 3: Performance of using different pre-trained language models: RoBERTa Nguyen et al. (2020) (top) and KEBERT Kawintiranon and Singh (2021) (bottom). The reported results are the Macro F1-score across all targets in a dataset on in-target multi-modal stance detection.
Figure 4: Performance of using different multi-modal fusion methods. The reported results are the Macro F1-score across all targets in a dataset on in-target multi-modal stance detection.
@@ 6.4 Generalization of Targeted Multi-modal Prompt Tuning
Previous experiments have demonstrated that the stance expression primarily resides in the textual modality. Therefore, to investigate the generalization of our TMPT when used with different pre-trained language models, we conduct experiments with two variants of our TMPT by using two other promising PLMs: RoBERTa Liu et al. (2019b) and KEBERT Kawintiranon and Singh (2021). The results are shown in Figure 3. Note that our Targeted Multi-modal Prompting can directly work with the two PLMs and achieve better performance, showing the compatibility of our method with various pre-trained models.
@@
We choose cross-modal attention Wei et al. (2020) and GAT Velickovic et al. (2018) to analyze the performance of our TMPT with different multi-modal fusion methods. The results are shown in Figure 4. It can be seen that no matter which fusion method we use, the experimental performance is better than BERT+ViT. This indicates that our TMPT can directly work with various multi-modal fusion methods and lead to improved multi-modal stance detection performance.
6.5 Error Analysis
From the results in Table 6, we can see that our TMPT performs well in the Mwtwt dataset, but still insufficient in other datasets on zero-shot stance detection. One possible reason is that targets in Mwtwt are all about expressing views on corporate mergers, so unknown targets can easily find commonalities in the dataset. However, for other datasets, the topics are diverse, which poses a challenge to mining targeted information. Therefore, how to better learn the correlation information between targets from data on diverse topics is a potential direction to improve the performance of the zero-shot scenario.
Further, after analyzing examples of misclassification, we found that among the incorrect samples, approximately 70% (based on calculations on randomly sampled 300 incorrect samples from five datasets) of the images contained stance-related content. This indicates that for our multi-modal stance detection task, it is important for further exploration in extracting and utilizing features from the visual modality.
Person Events Words Memes Mixed
ERR-Img(%) 7.8 13.4 37.7 26.6 14.5
ERR-Img-prop 0.28 0.60 1.60 1.95 1.16
Table 8: The results of error analysis on image types. ERR-Img means the proportion of different image types among the incorrect samples with stance-related content. ERR-Img-prop means ERR-Img divided by the proportion of its image types in the original dataset.
Building on the previous step, we conduct error analysis on different image types. The results are illustrated in Table 8. A larger ERR-Img-prop indicates a weaker ability of the model to handle samples with images of that category, indicating that for images containing words, memes, and mixed features, more effective methods of feature extraction and understanding remain to be proposed.
6.6 Visualization
@@ 7 Conclusion
In this paper, we present Mtse, Mccq, Mwtwt, Mruc and Mtwq, five new datasets for multi-modal stance detection. Based on the created datasets, we present in-target multi-modal stance detection and zero-shot multi-modal stance detection, aiming to advance and facilitate research in the field of multi-modal stance detection. In addition, we propose a simple yet effective targeted multi-modal prompting framework (TMPT) to deal with multi-modal stance detection, where the target information is explored to prompt the pre-trained models in learning multi-modal stance features. Extensive experiments on the new datasets show that our proposed TMPT achieves overall better performance than state-of-the-art baseline methods.
Acknowledgements
This work was partially supported by the National Natural Science Foundation of China (62176076), Natural Science Foundation of GuangDong 2023A1515012922, the Shenzhen Foundational Research Funding (JCYJ20220818102415032, JCYJ20210324115614039), the Major Key Project of PCL2021A06, Guangdong Provincial Key Labo-ratory of Novel Security Intelligence Technologies 2022B1212010005.
@@ A.1 Multi-modal Twitter Stance Election 2020
    DT 01/01/2020 $\rightarrow$ 09/30/2020 JB 01/01/2020 $\rightarrow$ 09/30/2020
@@ A.4 Multi-modal Russo-Ukrainian Conflict
    RUS 01/01/2022 $\rightarrow$ 06/30/2023 UKR 01/01/2022 $\rightarrow$ 06/30/2023
@@ A.5 Multi-modal Taiwan Question
    MOC 01/01/2022 $\rightarrow$ 06/30/2023 TOC 01/01/2022 $\rightarrow$ 06/30/2023
@@ C.1 Analysis of Textual Prompt Tuning
For textual prompt tuning, we utilize a frozen paradigm to make full use of the semantic information learned by the pre-trained language model. To analyze the effectiveness of the frozen paradigm, we also tried a tuned soft prompt, the results are shown in Table 9. The results of the tuned paradigm are extremely poorer than the frozen one and fluctuate greatly. One possible reason is that, unlike the continuous picture patch in the visual transformer, the token in the textual modality is discrete single words, so it is hard to use a gradient-based method to fine-tune the soft prompts like a visual modality. Thus, in this paper, we choose a manually designed fixed prompt to obtain better performance.
@@
Table 9: Experimental results of fixed prompt and tuned soft prompt of textual modality. The reported results are the Macro F1-score across all targets in a dataset on in-target multi-modal stance detection.
@@
To analyze the impact of the type of the hand-design textual prompts, we design several types of textual prompts and report the experiments in Table 10. Take the Mtse dataset as an example, it can be seen that the experimental results of all types of prompts are superior to those without prompts (BERT+ViT). This demonstrates the significance of the targeted prompts in this task. Further, we can also see that there are considerable differences in performance between different types of prompts. Specifically, when using both target and stance to design prompts (③, ④, and ⑤), the model performs significantly better. Therefore, in our method, we choose the type of ⑤ to devise the textual prompts for multi-modal stance detection.
@@
BERT+ViT - 53.29
Trump 58.54
Donald Trump 59.53
stance on Donald Trump 60.24
What is the stance on Donald Trump? 58.85
@@ C.2 Analysis of Visual Prompt Tuning
Following Jia et al. (2022), we conduct experiments with different depths of visual prompt tuning to analyze the impact of the depth of prompt tuning in our model. Here, shallow prompt tuning refers to only fine-tuning the prompt tokens in the Embedding layer of ViT, while deep prompt tuning refers to fine-tuning the prompt tokens in the Embedding layer of ViT and each layer in the transformer. The experimental results are shown in Table 11. We can see that shallow prompt tuning has a better overall effect.
Method Mtse Mccq Mwtwt Mruc Mtwq
Shallow (Ours) 60.84 67.67 77.59 75.76 67.59
@@
Table 11: Comparison results of using the shallow prompt tuning and deep prompt tuning. Shallow prompt tuning refers to only fine-tuning the prompt tokens in the Embedding layer of ViT, while deep prompt tuning refers to fine-tuning the prompt tokens in the Embedding layer of ViT and each layer in the Transformer. The reported results are the Macro F1-score across all targets in a dataset on in-target multi-modal stance detection.
@@
To analyze the impact of the number of visual prompt tokens, we set the value range of tokens as {3, 5, 7, 9} for comparative experiments. The results are shown in Figure 6. Note that different values of tokens can have a certain impact on performance. When the value is 7, the overall performance of the model on all datasets is the best. Therefore, we set the number of tokens to 7 in our method.
Figure 6: Experimental results of using different numbers of visual prompting tuning tokens. The reported results are the Macro F1-score across all targets in a dataset on in-target multi-modal stance detection.